Find the slope of the line through the points (1/2,3/4) and (3/5, 5/4)
Answer by srisanj4(2)

Slope of a line passing through two points (x1, y1) and (x2, y2) is given by
Slope m = (y2-y1)/(x2-x1)
Here, x1 = 1/2 y1 = 3/4
x2 = 3/5 y2 = 5/4
plugging in these values for x1, y1, x2 and y2
m = (5/4-3/4) / (3/5 - 1/2)
= 2/4 / 1/10
= 2 * 10/4*1
= 5
Hence the slope of line pass through said points is 5